OpenStudy (anonymous):

Multiply 1 over 6 multiplied by negative 1 over 2. negative 1 over 3 negative 1 over 12 1 over 12 1 over 3 HELP

OpenStudy (anonymous):

@Abhisar @Brainybeauty @Cydney_Morgan @dan815 @eric_d @ganeshie8 @HaileyD @insa @just_one_last_goodbye @Kainui @lolabieber @MrNood @nincompoop @ophercule @Preetha @quickstudent @RedNeckOutLaw @sbc4242 @taylordanielle @xoxoimher @YanaSidlinskiy @ziqbal103

OpenStudy (anonymous):

1/6 * (-1/2) * (-1/3) * (-1/12) * (1/12) * 1/3)

OpenStudy (anonymous):

can you tell me what the answer is please @Brainybeauty

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

multiply all the numbers in the denominator and you will get (-15552)

OpenStudy (anonymous):

so the answer is -1/15552

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea but that is not one of the answer choices @Brainybeauty

OpenStudy (anonymous):

these are the choices @brainybeauty negative 1 over 3 negative 1 over 12 1 over 12 1 over 3

OpenStudy (anonymous):

@Abhisar @Cydney_Morgan @dan815 @eric_d @ganeshie8 @HaileyD @insa @just_one_last_goodbye @Kainui @lolabieber @MrNood @nincompoop @ophercule @Preetha @quickstudent @RedNeckOutLaw @sbc4242 @taylordanielle @xoxoimher @YanaSidlinskiy @ziqbal103

OpenStudy (anonymous):

\(\dfrac{1}{6}-\dfrac{1}{2}-\dfrac{1}{3}-\dfrac{1}{12}*\dfrac{1}{12}*\dfrac{1}{3}\) is this ur question?